OpenStudy (anonymous):

heeelpp solve using substitution method. y=x=3 y=5x-5

OpenStudy (anonymous):

y = x - 3 -x = -y - 3 /-1 x = y + 3 substitute that into the other equation: y=5x-5 y = 5(y + 3) - 5 y = 5y + 15 - 5 combine like terms: -4y = 10 y = -2.5 Now put the y-value into one of the original equations and solve for x.

OpenStudy (anonymous):

Hope this helps:)